OpenStudy (anonymous):

A wooden pyramid, 6 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 3 and makes a larger pyramid with the new dimensions. Describe the ratio of the volumes of the two pyramids.

OpenStudy (anonymous):

please give me a min to solve

OpenStudy (anonymous):

Ok; appreciate your time.

OpenStudy (anonymous):

ratio of measurements is 1:3 ratio of volumes (1)^3 :(3)^3 =1:27 ok suppose that h= 6 and the square base has side length= 1cm V=1/3×12×6 V = 2 cm^3 now with the dimensions tripled h = 18 and square side length = 3cm V=1/3×32×18 V = 54 cm^3 the ratio of the volumes is 2:54 or 1: 27 hope this will work :)

OpenStudy (anonymous):

Yes it does, thank you!

OpenStudy (anonymous):

welcome :)